What would be the yield to call if the call can be made in 3 years at a price of $1,100 on a bond currently priced at $1,050 with a 5% coupon rate and $1,000 par value?

A. 6.23% B. 6.25% C. 3.24% D. 3.22%
Answer #1

CASE 1: ASSUMING ANNUAL COUPON PAYMENTS
Using financial calculator
N=3
PMT=5%*1000
PV=-1050
FV=1000
CPT I/Y=3.22%

Hence, yield to call=3.22%

CASE 2: ASSUMING SEMI-ANNUAL COUPON PAYMENTS
Using financial calculator
N=3*2
PMT=5%*1000/2
PV=-1050
FV=1000
CPT I/Y=1.6188%

Hence, yield to call=3.24%